town B is 1124 km north of city A, an airplane that can travel at 765 km/h in still air departs city A and plans to fly in a straight line to city B, despite wind blowing at 65.6 km/h W. The magnitude of the planes velocity relative to the ground is?


1
Expert's answer
2021-11-21T17:24:52-0500

"v=\sqrt{765^2-65.6^2}=762\ (km/h)"


"\tan\theta=\frac{65.6}{762}=0.086\to \theta\approx5°\ W\ of \ N"